# From Prompt Injections to Protocol Exploits: Threats in LLM-Powered AI Agents Workflows

### v1

@@ VI-A Model‐Control‐Protocol (MCP)
MCP [58] and A2A [59] are complementary in enabling robust communication among LLM-powered agents. MCP (Model Context Protocol) is a standardized bridge between a language model and the external resources it requires during inference. By defining a client–server architecture in which MCP clients embedded within agent hosts request specific capabilities—such as file retrieval, database queries, or external API calls from lightweight MCP servers—agents can dynamically augment their context window without hardcoding tool integrations. In this way, MCP enables vertical integration, allowing an LLM to invoke structured tools and access relevant information (for example, fetching a document or executing a function) in real-time, which enhances the agent’s reasoning with up-to-date, application-specific data.
@@
  • Credential theft: token theft/replay, confused-deputy proxy abuse, and insecure configuration. Attackers steal or misuse authentication tokens to impersonate agents, hijacking privileged MCP operations.
VI-B Agent2Agent (A2A) protocol
A2A (Agent-to-Agent) [59] extends this framework horizontally by orchestrating secure, task-oriented interactions among multiple autonomous agents. Leveraging a JSON-RPC-over-HTTP(S) communication model with Server-Sent Events for streaming updates, A2A enables one agent (the client) to discover and delegate subtasks to another agent (the remote) based on metadata published in “Agent Cards.” Each task is a defined object whose lifecycle—from initiation to artifact delivery—is governed by the protocol, allowing agents to coordinate complex workflows and negotiate content formats (text, images, forms, etc.) without sharing internal memory or tools. When combined, MCP equips each agent with the precise context it needs, while A2A orchestrates multi-agent collaboration, supporting scalable, interoperable ecosystems of LLM-enabled agents [62].
@@ VI-B Agent2Agent (A2A) protocol
  • Agent discovery spoofing: forged discovery responses or Agent-Card spoofing to impersonate malicious agents. Attackers present fake capability metadata to divert tasks to rogue agents under their control.
@@
  • Memory subversion: memory-poisoning (MINJA) within the agent framework. Injected memory artifacts corrupt internal state, causing unpredictable or malicious agent behavior.
Tab. VI summarizes the attack vectors across MCP [58], A2A[59], ANP[60] & ACP[61] protocols.
@@ VII Open Challenges and Future Directions
As LLMs and agentic systems become increasingly pervasive, securing these complex systems against sophisticated threats remains a significant challenge. This section identifies the key areas where current defenses fall short and outlines promising directions for future research.
@@ VII-A MCP protocol vulnerabilities
The Model Context Protocol (MCP) provides a unified, plug-and-play framework—akin to a “USB-C port for AI” for connecting large language models to diverse data sources and tool endpoints. By decoupling responsibilities into hosts (applications), clients (connectors), and servers (modular adapters), MCP greatly accelerates the assembly of complex, LLM-driven workflows. Yet this very modularity widens the attack surface: every integration point—from prompt handling to long-running connections—can be exploited if left unchecked [63].
The first significant challenge lies in establishing dynamic trust and adaptive policy enforcement within MCP deployments. Traditional static access controls cannot keep pace with the fluid, language-driven workflows that modern agents require. Future research should explore declarative policy languages that allow hosts to negotiate and update least-privilege rules at runtime, as well as distributed trust-negotiation protocols capable of granting time-boxed permissions for specific tool invocations without human intervention [25].
A second frontier concerns the integrity of context and provenance tracking. Because MCP maintains conversational state across multiple turns and service calls, any undetected tampering can subtly bias downstream actions. Promising directions include using secure enclaves or hardware root-of-trust modules to vault critical session memory and immutable audit logs—perhaps built on lightweight blockchain ledgers—to attest to the sequence of context updates cryptographically [64].
Ransomware[65] can cripple an MCP deployment by encrypting or locking access to the very data sources and tool-endpoints that MCP servers expose, effectively severing the “USB-C port” connections between LLMs and critical context (e.g., local files, databases, or APIs). Because MCP’s client–server architecture allows a host to plug into multiple lightweight servers, a single compromised server can cascade, denying clients the ability to fetch prompts, workflows, or sampling configurations and halting AI-driven operations until a ransom is paid directly undermining MCP’s goals of flexible, secure integration within your infrastructure.
@@ VII-B Agentic Web Interfaces: Design and Security Challenges
Lu et al. [66] present a forward-thinking perspective on advancing web-based AI agents by proposing the concept of an Agentic Web Interface (AWI), a new interaction paradigm tailored to the unique capabilities and limitations of LLM-powered agents. As LLMs and their multimodal counterparts increasingly demonstrate potential in automating complex web tasks, the fundamental misalignment between human-centric web interfaces and agent requirements has become a critical bottleneck. Current approaches, which attempt to adapt agents to navigate intricate web inputs such as dense Document Object Model (DOM) trees, screenshot-based environments, or direct API interactions, have proven insufficient in addressing these challenges. The authors introduce AWI, a standardized and agent-optimized interface that facilitates more efficient, reliable, and transparent interactions between AI agents and web environments, thereby overcoming these limitations. They outline six core design principles for AWI, focusing on safety, efficiency, and standardization, and emphasize the importance of collaboration across the machine learning community to realize this vision.
Despite its promise, the introduction of AWI also raises several open problems and security-related challenges that must be addressed to ensure safe deployment. First, introducing agent-optimized interfaces may inadvertently create new vulnerabilities, including unauthorized agent access, manipulation of AWI components, or exploitation of poorly implemented standardization protocols. Furthermore, as AWI enables more autonomous agent behaviors, attackers may exploit LLM-driven agents for malicious activities such as automated phishing, data exfiltration, or denial-of-service attacks within web environments. Future research should focus on developing robust security mechanisms for AWI, including authentication, integrity verification, and real-time anomaly detection, to monitor agent behavior effectively. Additionally, establishing secure communication standards and incorporating adversarial robustness testing for LLM-agent interactions will be essential. Promising research directions include exploring formal verification techniques for AWI protocols, creating secure sandbox environments for agent testing, and developing mechanisms for transparent auditing of agent actions. Addressing these vulnerabilities is crucial to ensure that the shift toward agent-optimized web interactions enhances functionality and security in LLM-powered agent ecosystems.
@@ VII-C Optimizing LLM Multi-Agent Systems: Design and Security
Zhou et al. [67] investigate the design complexity of Large Language Model (LLM)-powered multi-agent systems (MAS) and introduce an automated optimization framework, Multi-Agent System Search (MASS), to streamline the process. Recognizing that prompts and interaction topologies are critical for enabling effective collaboration among agents, the authors conduct a comprehensive analysis of the MAS design space. Building on these insights, they present MASS, which employs a staged optimization process that iteratively refines prompts and system topologies across three levels: block-level prompt optimization, workflow topology optimization, and global prompt optimization. Experimental results demonstrate that systems designed with MASS significantly outperform existing MAS approaches. The authors conclude by outlining key design principles for building robust, high-performing multi-agent systems.
Despite these advancements, several open challenges and vulnerabilities remain unaddressed in the design and deployment of LLM-driven MAS. The automation of prompt and topology generation, while improving system efficiency, introduces new risks such as the emergence of unpredictable agent behaviors, prompt injection attacks, and adversarial manipulation of communication topologies. Moreover, as MAS grow more complex, detecting unintended coordination patterns or hidden vulnerabilities becomes increasingly difficult. Future research should focus on integrating security constraints directly into the MASS optimization process, alongside developing verification techniques to analyze the safety and robustness of optimized multi-agent configurations. Promising directions also include designing resilient topologies that can withstand adversarial disruptions, incorporating defensive prompting strategies, and exploring real-time monitoring tools to detect and mitigate malicious agent interactions in LLM-powered MAS.
@@ VII-D VLM-Powered Web Agents and Security Challenges
Andreux et al. [68] introduce Surfer-H, a cost-efficient web agent that leverages Vision-Language Models (VLMs) to perform user-defined tasks within web environments autonomously. Central to this system is Holo1, an open-weight VLM collection specifically trained for web navigation and information extraction. Holo1 is developed using a diverse dataset that combines open-access web content, synthetic examples, and self-generated agentic data. Experimental results demonstrate that Holo1 outperforms general-purpose UI benchmarks and achieves leading results on the newly proposed WebClick benchmark for web UI localization. When integrated with Holo1, Surfer-H achieves a state-of-the-art accuracy of 92.2% on the WebVoyager benchmark, offering a strong balance between task performance and cost efficiency. Therefore, several open challenges and security vulnerabilities arise with the use of VLM-powered web agents, such as Surfer-H. Integrating vision and language capabilities for autonomous web interaction raises concerns related to interface manipulation, adversarial attacks on visual perception, and misuse of automated browsing for malicious purposes such as data scraping or phishing. Furthermore, the reliance on agent-generated data during training may unintentionally introduce biased behaviors or security flaws. Future research should focus on strengthening the robustness of VLM-based web agents against adversarial inputs, ensuring secure interaction with web environments, and establishing guidelines to prevent the abuse of agentic web technologies. Efforts toward explainability, real-time anomaly detection, and secure agent deployment practices will be essential for building reliable and trustworthy web agent ecosystems.
@@ VII-E Memory-Centric LLMs for Long-Term Dialogue and Emerging Security Risks
Chhikara et al. [69] present Mem0, a scalable memory-centric architecture designed to overcome the inherent limitations of Large Language Models (LLMs) in maintaining consistency over long, multi-session dialogues. Recognizing the challenges posed by fixed context windows, Mem0 dynamically extracts, consolidates, and retrieves relevant information from conversations to ensure contextual coherence. The authors also introduce an enhanced variant that incorporates graph-based memory structures to model complex relationships among conversational elements. Through extensive evaluations on the LOCOMO benchmark, the proposed approaches are compared against six established memory and retrieval baselines, including memory-augmented systems, RAG techniques, full-context models, and proprietary solutions. The results demonstrate that both Mem0 and its graph-based extension outperform all competing methods across diverse question categories, while significantly reducing computational overhead. Mem0 achieves notable improvements in accuracy and efficiency, highlighting the value of persistent, structured memory for advancing long-term conversational capabilities in LLM-driven agents.
While Mem0 advances the state of memory-enhanced LLMs, its reliance on structured memory mechanisms also introduces new vulnerabilities that require attention. The processes of information extraction, consolidation, and retrieval within long-term memory could be exploited by adversaries through data poisoning, manipulation of stored knowledge, or injection of misleading conversational elements. Additionally, the use of graph-based memory structures raises concerns about unauthorized inference of hidden relationships or the leakage of sensitive information. Future research should focus on securing memory retrieval pipelines, implementing tamper-resistant memory architectures, and developing mechanisms for real-time monitoring of memory integrity to ensure robustness and security. Promising directions include integrating privacy-preserving techniques, adversarial robustness testing for memory components, and formal verification of memory consistency to ensure both the effectiveness and security of LLM-powered conversational agents.

### v2

@@ VI-B Model‐Control‐Protocol (MCP)
MCP [modelcontextprotocol2025] and A2A [googleblog2025] are complementary in enabling robust communication among LLM-powered agents. MCP (Model Context Protocol) is a standardized bridge between a language model and the external resources it requires during inference. By defining a client–server architecture in which MCP clients embedded within agent hosts request specific capabilities—such as file retrieval, database queries, or external API calls — from lightweight MCP servers, agents can dynamically augment their context window without hardcoding tool integrations. In this way, MCP enables vertical integration, allowing an LLM to invoke structured tools and access relevant information (e.g., fetching a document or executing a function) in real time, thereby enhancing the agent’s reasoning with up-to-date, application-specific data.
@@
  • Credential theft: token theft/replay, confused-deputy proxy abuse, and insecure configuration. Attackers steal or misuse authentication tokens to impersonate agents and hijack privileged MCP operations.
Table VIII presents representative CVEs from 2025 cross-mapped to the four proposed threat model categories, demonstrating the empirical grounding and real-world relevance of our taxonomy. These vulnerabilities span indirect prompt injection (e.g., LangChain’s GmailToolkit), server-side request forgery, XML-based data exfiltration, path traversal, protocol misconfigurations (e.g., in MCP-Remote), and sandbox escapes in agent tools. Notably, most of these attacks exploit insecure interfaces, permissive plugin logic, or protocol-level flaws—often requiring only low privileges or indirect input access—highlighting the brittleness of current LLM-agent integrations. This validates the practical urgency of securing prompt surfaces, agent-tool bridges, and multi-agent coordination flows through structured, layered defenses.
VI-C Agent2Agent (A2A) protocol
A2A (Agent-to-Agent) [googleblog2025] extends this framework horizontally by orchestrating secure, task-oriented interactions among multiple autonomous agents. Leveraging a JSON-RPC-over-HTTP(S) communication model with Server-Sent Events for streaming updates, A2A enables one agent (the client) to discover and delegate subtasks to another agent (the remote) based on metadata published in “Agent Cards.” Each task is a defined object whose lifecycle—from initiation to artifact delivery—is governed by the protocol, allowing agents to coordinate complex workflows and negotiate content formats (text, images, forms, etc.) without sharing internal memory or tools. When combined, MCP equips each agent with the precise context it needs, while A2A orchestrates multi-agent collaboration, supporting scalable, interoperable ecosystems of LLM-enabled agents [habler2025building].
@@ VI-C Agent2Agent (A2A) protocol
  • Agent discovery spoofing: forged discovery responses or Agent-Card spoofing to impersonate malicious agents. The internal state leads to capability metadata to divert tasks to rogue agents under their control.
@@
  • Memory subversion: memory-poisoning (MINJA) within the agent framework. Injected memory artifacts corrupt the internal state, leading to unpredictable or malicious agent behavior.
Tab. VII summarizes the attack vectors across MCP [modelcontextprotocol2025], A2A[googleblog2025], ANP[agentnetworkprotocol2024] & ACP[agentcommunicationprotocol2024] protocols.
CVE ID Date Component Affected Summary of the Issue CWE-ID Threat Category
CVE-2025-46059 [CVE-2025-46059] 2025 LangChain – GmailToolkit (v0.3.51) Critical indirect prompt injection via crafted emails enabling arbitrary code execution. CWE-94 [CWE-94] Input Manipulation
CVE-2025-2828 [CVE-2025-2828] 2025 LangChain – RequestsToolkit (OpenAPI tool) SSRF flaw bypasses URL filters, allowing internal network access. CWE-918 [CWE-918] Protocol Vulnerabilities
CVE-2025-6984 [CVE-2025-6984] 2025 LangChain – EverNoteLoader (v0.3.63) XML External Entity (XXE) attacks via insecure XML parsing expose sensitive files. CWE-200 [CWE-200] System & Privacy Attacks
CVE-2025-6985 [CVE-2025-6985] 2025 LangChain – HTMLSectionSplitter (v0.3.8) Unsafe XSLT parsing allows file disclosure via malicious stylesheets. CWE-611 [CWE-611] System & Privacy Attacks
CVE-2025-6853 [CVE-2025-6853] 2025 LangChain-Chatchat (v0.3.1) – Backend API Path traversal in temp upload allows out-of-scope file creation. CWE-22 [CWE-22] System & Privacy Attacks
CVE-2025-6854 [CVE-2025-6854] 2025 LangChain-Chatchat (v0.3.1) – Backend API Base64 path traversal lets an attacker read unintended files. CWE-22 [CWE-22] System & Privacy Attacks
CVE-2025-6855 [CVE-2025-6855] 2025 LangChain-Chatchat (v0.3.1) – Backend API Path traversal in file uploads allows overwriting critical files. CWE-22 [CWE-22] System & Privacy Attacks
CVE-2025-11844 [CVE-2025-11844] 2025 Hugging Face – SmolAgents (v1.20.0) XPath injection lets attacker bypass filters, access unauthorized DOM elements. CWE-643 [CWE-643] Input Manipulation
CVE-2025-59532 [CVE-2025-59532] 2025 OpenAI – Codex CLI (v0.2.0–0.38.0) Sandbox escape via directory control flaw allows arbitrary file execution. CWE-20 [CWE-20] Model Compromise
CVE-2025-6514 [CVE-2025-6514] 2025 Anthropic – MCP-Remote (v0.0.5–0.1.15) Untrusted MCP endpoints can trigger remote OS command execution. CWE-78 [CWE-78] Protocol Vulnerabilities
CVE-2025-49596 [CVE-2025-49596] 2025 Anthropic – MCP Inspector (v0.14.0) CSRF/DNS-rebinding vulnerability enables RCE via Inspector UI. CWE-306 [CWE-306] Protocol Vulnerabilities
TABLE VIII: Representative CVEs cross-mapped to the four threat model categories. CWE-94 (Code Injection), CWE-918 (Server-Side Request Forgery), CWE-200 (Information Exposure), CWE-611 (Improper Restriction of XML External Entity Reference), CWE-22 (Path Traversal), CWE-643 (XPath Injection), CWE-20 (Improper Input Validation), CWE-78 (OS Command Injection), and CWE-306 (Missing Authentication for Critical Function).
@@ VII Open Challenges and Future Directions
As LLMs and agentic systems become increasingly pervasive, securing them against sophisticated threats remains a significant challenge. This section identifies the key areas where current defenses fall short and outlines promising directions for future research.
@@ VII-A MCP protocol vulnerabilities
The Model Context Protocol (MCP) provides a unified, plug-and-play framework—akin to a “USB-C port for AI” for connecting large language models to diverse data sources and tool endpoints. By decoupling responsibilities into hosts (applications), clients (connectors), and servers (modular adapters), MCP greatly accelerates the assembly of complex, LLM-driven workflows. Yet this very modularity widens the attack surface: every integration point—from prompt handling to long-running connections—can be exploited if left unchecked [hasan2025model].
The first significant challenge lies in establishing dynamic trust and adaptive policy enforcement within MCP deployments. Traditional static access controls cannot keep pace with the fluid, language-driven workflows that modern agents require. Future research should explore declarative policy languages that allow hosts to negotiate and update least-privilege rules at runtime, as well as distributed trust-negotiation protocols capable of granting time-boxed permissions for specific tool invocations without human intervention [hou2025model].
A second frontier concerns the integrity of context and provenance tracking. Because MCP maintains conversational state across multiple turns and service calls, any undetected tampering can subtly bias downstream actions. Promising directions include using secure enclaves or hardware root-of-trust modules to vault critical session memory and immutable audit logs—perhaps built on lightweight blockchain ledgers—to attest to the sequence of context updates cryptographically [ahmadi2025mcp].
Ransomware[al2024securing] can cripple an MCP deployment by encrypting or locking access to the very data sources and tool-endpoints that MCP servers expose, effectively severing the “USB-C port” connections between LLMs and critical context (e.g., local files, databases, or APIs). Because MCP’s client–server architecture allows a host to plug into multiple lightweight servers, a single compromised server can cascade, denying clients the ability to fetch prompts, workflows, or sampling configurations and halting AI-driven operations until a ransom is paid, directly undermining MCP’s goals of flexible, secure integration within your infrastructure.
@@ VII-B Agentic Web Interfaces: Design and Security Challenges
Lu et al. [lu2025build] present a forward-thinking perspective on advancing web-based AI agents by proposing the Agentic Web Interface (AWI), a new interaction paradigm tailored to the unique capabilities and limitations of LLM-powered agents. As LLMs and their multimodal counterparts increasingly demonstrate the potential to automate complex web tasks, the fundamental misalignment between human-centric web interfaces and agent requirements has become a critical bottleneck. Current approaches, which attempt to adapt agents to navigate intricate web inputs such as dense Document Object Model (DOM) trees, screenshot-based environments, or direct API interactions, have proven insufficient in addressing these challenges. The authors introduce AWI, a standardized, agent-optimized interface that enables more efficient, reliable, and transparent interactions between AI agents and web environments, thereby overcoming these limitations. They outline six core design principles for AWI, focusing on safety, efficiency, and standardization, and emphasize the importance of collaboration across the machine learning community to realize this vision.
Despite its promise, the introduction of AWI also raises several open problems and security-related challenges that must be addressed to ensure safe deployment. First, introducing agent-optimized interfaces may inadvertently create new vulnerabilities, including unauthorized agent access, manipulation of AWI components, or exploitation of poorly implemented standardization protocols. Furthermore, as AWI enables more autonomous agent behaviors, attackers may exploit LLM-driven agents for malicious activities such as automated phishing, data exfiltration, or denial-of-service attacks within web environments. Future research should focus on developing robust security mechanisms for AWI, including authentication, integrity verification, and real-time anomaly detection, to effectively monitor agent behavior. Additionally, establishing secure communication standards and incorporating adversarial robustness testing for LLM-agent interactions will be essential. Promising research directions include exploring formal verification techniques for AWI protocols, creating secure sandbox environments for agent testing, and developing mechanisms for transparent auditing of agent actions. Addressing these vulnerabilities is crucial to ensure that the shift toward agent-optimized web interactions enhances functionality and security in LLM-powered agent ecosystems.
@@ VII-C Optimizing LLM Multi-Agent Systems: Design and Security
Zhou et al. [zhou2025multi] investigate the design complexity of Large Language Model (LLM)-powered multi-agent systems (MAS) and introduce an automated optimization framework, Multi-Agent System Search (MASS), to streamline the process. Recognizing that prompts and interaction topologies are critical for enabling effective collaboration among agents, the authors conduct a comprehensive analysis of the MAS design space. Building on these insights, they present MASS, which employs a staged optimization process that iteratively refines prompts and system topologies across three levels: block-level prompt optimization, workflow topology optimization, and global prompt optimization. Experimental results demonstrate that systems designed with MASS significantly outperform existing MAS approaches. The authors conclude by outlining key design principles for building robust, high-performing multi-agent systems.
Despite these advancements, several open challenges and vulnerabilities remain unaddressed in the design and deployment of LLM-driven MAS. The automation of prompt and topology generation, while improving system efficiency, introduces new risks, including unpredictable agent behavior, prompt injection attacks, and adversarial manipulation of communication topologies. Moreover, as MAS grow more complex, detecting unintended coordination patterns or hidden vulnerabilities becomes increasingly difficult. Future research should focus on integrating security constraints directly into the MASS optimization process, alongside the development of verification techniques to analyze the safety and robustness of optimized multi-agent configurations. Promising directions also include designing resilient topologies that can withstand adversarial disruptions, incorporating defensive prompting strategies, and exploring real-time monitoring tools to detect and mitigate malicious agent interactions in LLM-powered MAS.
@@ VII-D VLM-Powered Web Agents and Security Challenges
Andreux et al. [andreux2025surfer] introduce Surfer-H, a cost-efficient web agent that leverages Vision-Language Models (VLMs) to perform user-defined tasks within web environments autonomously. Central to this system is Holo1, an open-weight VLM collection specifically trained for web navigation and information extraction. Holo1 is developed using a diverse dataset that combines open-access web content, synthetic examples, and self-generated agentic data. Experimental results demonstrate that Holo1 outperforms general-purpose UI benchmarks and achieves leading results on the newly proposed WebClick benchmark for web UI localization. When integrated with Holo1, Surfer-H achieves a state-of-the-art accuracy of 92.2% on the WebVoyager benchmark, offering a strong balance between task performance and cost efficiency. Therefore, several open challenges and security vulnerabilities arise when using VLM-powered web agents, such as Surfer-H. Integrating vision and language capabilities for autonomous web interaction raises concerns about interface manipulation, adversarial attacks on visual perception, and the misuse of automated browsing for malicious purposes, such as data scraping or phishing. Furthermore, the reliance on agent-generated data during training may unintentionally introduce biased behaviors or security flaws. Future research should focus on strengthening the robustness of VLM-based web agents against adversarial inputs, ensuring secure interaction with web environments, and establishing guidelines to prevent the abuse of agentic web technologies. Efforts toward explainability, real-time anomaly detection, and secure agent deployment practices will be essential for building reliable and trustworthy web agent ecosystems.
@@ VII-E Memory-Centric LLMs for Long-Term Dialogue and Emerging Security Risks
Chhikara et al. [chhikara2025mem0] present Mem0, a scalable memory-centric architecture designed to overcome the inherent limitations of Large Language Models (LLMs) in maintaining consistency over long, multi-session dialogues. Recognizing the challenges posed by fixed context windows, Mem0 dynamically extracts, consolidates, and retrieves relevant information from conversations to ensure contextual coherence. The authors also introduce an enhanced variant that incorporates graph-based memory structures to model complex relationships among conversational elements. Through extensive evaluations on the LOCOMO benchmark, the proposed approaches are compared against six established memory and retrieval baselines, including memory-augmented systems, RAG techniques, full-context models, and proprietary solutions. The results demonstrate that both Mem0 and its graph-based extension outperform all competing methods across diverse question categories, while significantly reducing computational overhead. Mem0 achieves notable improvements in accuracy and efficiency, highlighting the value of persistent, structured memory for advancing long-term conversational capabilities in LLM-driven agents.
While Mem0 advances the state of memory-enhanced LLMs, its reliance on structured memory mechanisms also introduces new vulnerabilities that require attention. The processes of information extraction, consolidation, and retrieval within long-term memory could be exploited by adversaries through data poisoning, manipulation of stored knowledge, or injection of misleading conversational elements. Additionally, the use of graph-based memory structures raises concerns about unauthorized inference of hidden relationships or the leakage of sensitive information. Future research should focus on securing memory retrieval pipelines, implementing tamper-resistant memory architectures, and developing mechanisms for real-time monitoring of memory integrity to ensure robustness and security. Promising directions include integrating privacy-preserving techniques, testing for adversarial robustness in memory components, and formal verification of memory consistency to ensure the effectiveness and security of LLM-powered conversational agents.